OpenStudy (anonymous):

show that the conjecture is false by finding a counter example. For any real number x, x^3 > or equal to x^2. is x as 1/4 a liable answer?

OpenStudy (anonymous):

Well, you can check! \[ \left( \frac{1}{4} \right)^3 = \frac{1}{64} \] \[ \left( \frac{1}{4} \right)^2 = \frac{1}{16} \] and \[ \frac{1}{64} < \frac{1}{16} \] So you're correct :)